# Contextual Knowledge Graph

### v2

@@ Abstract
Knowledge Graphs (KGs) are foundational structures in many AI applications, representing entities and their interrelations through triples. However, triple-based KGs lack the contextual information of relational knowledge, like temporal dynamics and provenance details, which are crucial for comprehensive knowledge representation and effective reasoning. Instead, Contextual Knowledge Graphs (CKGs) expand upon the conventional structure by incorporating additional information such as time validity, geographic location, and source provenance. This integration provides a more nuanced and accurate understanding of knowledge, enabling KGs to offer richer insights and support more sophisticated reasoning processes. In this work, we first discuss the inherent limitations of triple-based KGs and introduce the concept of contextual KGs, highlighting their advantages in knowledge representation and reasoning. We then present KGR3, a context-enriched KG reasoning paradigm that leverages large language models (LLMs) to retrieve candidate entities and related contexts, rank them based on the retrieved information, and reason whether sufficient information has been obtained to answer a query. Our experimental results demonstrate that KGR3 significantly improves performance on KG completion (KGC) and KG question answering (KGQA) tasks, validating the effectiveness of incorporating contextual information on KG representation and reasoning.
Contextual Knowledge Graph
@@ 1 Introduction
While this triple-based structure offers clear semantics and precision through the use of schemas and ontologies, it loses the contextual information of knowledge and falls short in capturing the complexity and richness of real-world knowledge Dong (2023). Since we cannot clearly model the knowledge in a domain only with entities and relations, many recent KGs Pellissier Tanon et al. (2020); Tharani (2021) are designed to be semi-structured: they leverage the clear semantics of structured data provided by the rigidity of schemas (i.e., ontologies) while also embracing the flexibility of unstructured data. Such KGs integrate multi-modal knowledge, including entity description, images, timestamps and other metadata, all of which can be regarded as the contexts of triple knowledge. In this paper, we refer to this type of KGs as contextual KGs (CKGs). By incorporating these semantic contexts, contextual KGs provide a more comprehensive and nuanced representation of knowledge, extending beyond the traditional triple-based approach. This enables KGs to possess more advanced capabilities in knowledge representation and reasoning.
Moreover, large language models (LLMs), pre-trained on vast text corpora, have exhibited strong semantic understanding capability Brown et al. (2020a). And the use of LLMs for KG reasoning has become a research hotspot Wei et al. (2023); Liao et al. (2024); Sun et al. (2024). However, KGs may contain numerous entities and relations, but not all entities and relationships are fully annotated and connected, leading to data sparsity. This sparsity results in a lack of sufficient contextual information for the LLM during inference. On the other hand, LLMs are better at handling unstructured data rather than structured triples. Considering that contextual KGs can provide unstructured contextual information for LLM reasoning, the synergy between contextual KGs and LLMs holds significant potential for advancing the field of knowledge reasoning.
In this paper, we will first give a brief discussion on the limitations of the triple-based KGs and give the specific definition of contextual KGs. To validate the effectiveness of contextual KG on enhancing KG representation and reasoning, we propose a novel context-enhanced KG reasoning paradigm, named KGR3, which leverage the strong reasoning power of LLMs to firstly retrieve candidate entities and related contexts from KG, and rank the candidate entities based on retrieved context, and then reason whether sufficient information is retrieved to answer the question. Experimental results demonstrate that our proposed paradigm KGR3 enhances the performance of existing models on the tasks of KG completion (KGC) and KG question answering (KGQA), which are two of fundamental reasoning tasks over KGs. Experimental
@@
  • Point out the limitations of the current triple-based KGs, and introduce the concept of contextual KG, which has more advanced capabilities in knowledge representation and reasoning.
  • Propose a context-enhanced KG reasoning paradigm, KGR3, which leverages the LLM to perform KG reasoning based on related contexts. Experimental results on KGC and KGQA support our intuition that the integration of contextual data can contribute to effective KG reasoning.
2 Contextual Knowledge Graph
In this section, we first discuss on the limitations of triple-based KGs, caused by the absence of contextual information. Moreover, we point out the effects of contextual information on knowledge representation and reasoning, then categorize and interpreter different types of contexts in the KGs. Finally, we formally define contextual KGs as well as two reasoning tasks over contextual KGs, i.e. KG completion (KGC) and KG question answering (KGQA).
@@ 2.2 The Effects of Contextual Information
To address the limitations of triple-based KGs, a promising approach is to attach contextual data to factual triples. For instance, several KGs, such as YAGO and the Yahoo Knowledge Graph, include meta-information with their facts, such as the time of validity, the geographic location of a fact, and provenance information. By integrating such data, contextual KGs can offer a more comprehensive and accurate representation of knowledge, thereby enabling more effective reasoning.
@@
During the process of knowledge reasoning, contextual information within contextual KGs can be leveraged to associate entities that are not directly connected by identifying similar contexts. This capability is particularly useful for making connections and drawing inferences that go beyond the predefined relation set of a triple-based KG. Moreover, contextual information provides additional knowledge, allowing for larger knowledge coverage and greater flexibility compared to triple-based KGs. Specifically, contextual information can be used to answer complex reasoning questions, such as those involving qualifiers or specific conditions that are often hidden within contextual data. For instance, answering a question about "which company is Apple’s biggest competitor in the global smartphone market" would require integrating quantitative data, temporal information, and detailed market dynamics analysis with basic entity and relation information in KGs, as shown in Figure 1(d). Contextual KGs thus enable the handling of such intricate queries by providing a richer and more detailed knowledge base.
2.3 Categories of Contextual Data in KGs
@@
As shown in Figure 2, contextual data in KGs can be roughly classified into two categories, i.e., entity contexts and triple contexts.
@@ 2.4 Problem Specification
A Contextual Knowledge Graph (denoted as $\mathcal{CKG}=\{\mathcal{E},\mathcal{R},\mathcal{Q},\mathcal{EC},\mathcal{RC}\}$) can be represented as a set of factual quadruples in the form of $(h,r,t,rc)\in\mathcal{Q}$, where $h,t\in\mathcal{E}$, $r\in\mathcal{R}$ and $rc\in\mathcal{RC}$. The notations $h$ and $t$ denote the head and the tail entity of a factual quadruple, $r$ denotes the relations between $h$ and $t$, and $rc$ denotes . $\mathcal{EC},\mathcal{RC}$ are the set of entity contexts and relation contexts. Each entity $e\in\mathcal{E}$ and its entity context $ec\in\mathcal{EC}$ form a complete entity representation $(e,ec)$.
@@
Given a natural question $nq$ and its topic entity $e_{topic}\in\mathcal{E}$, KGQA aims to retrieve related knowledge by generating structured queries or sampling subgraphs from $\mathcal{KG}$ and predict the answer $a$ based on retrieved knowledge, i.e, $a=f(nq,e_{topic},\mathcal{KG})$. For performing QA over a contextual KG, the prediction function can be reformulated as $f(nq,e_{topic},\mathcal{CKG})$.
Figure 3: The pipeline of the KGR3 paradigm.
@@ 3 Methods
In this section, we introduce KGR3, a novel KG reasoning paradigm that leverages LLMs to perform KG reasoning tasks based on tructured and contextual semantics. We aim to utilize the complementary relationship between both semantics to improve the reliability and explainability of the reasoning process.
As shown in For triple-based KGs, we begin by augmenting the KG with necessary contextual information extracted from relevant databases, a step that can be omitted if the KG is already a contextual KG. The KGR3 paradigm consists of three main steps: The Retrieval step is to retrieve candidate entities and related contexts from the contextual KG based on the given question; the Ranking step involves ranking candidate entities based on the contexts and the given question; the Reasoning step is to exploit the LLM to determine whether sufficient information is retrieved. If sufficient information is available, the answer will be generated. If not, the whole processes iterates by retrieving new information based on the top-ranked candidate entities. We give a detailed description of the proposed context-aware paradigms for the KGC and the KBQA tasks
@@ 3.1 Context Extraction
Currently, commonly used KG datasets, such as FB15k237, YAGO3-10, and Wikidata5M, are encyclopedic KGs that encapsulate general knowledge about the real world. These KGs are typically developed by domain experts by applying named entity recognition and relation extraction techniques on Wikipedia. However, during this construction process, the rich contexts surrounding the entities are often omitted. Recent studies Wang et al. (2021, 2022b) have proposed to incorporate entity labels and descriptions as supplementary information for KGs. Nevertheless, the labels and descriptions are insufficient to replace the specific contexts associated with KG triples, thereby limiting their effectiveness in addressing diverse KG reasoning problems.
@@ 3.2 Knowledge Graph Completion
In this section, we demonstrate a new context-enriched KGC method based on our proposed KGR3 paradigm. Since KGC can be considered as an entity ranking task for single-hop reasoning questions, it is not necessary to perform iterative reasoning processes. Thus, the reasoning step is omitted for this task.
@@ 3.3.3 Step 3: Context-aware Reasonin
Upon obtaining the new top-$M$ reasoning paths $P$ and updating relation context list $C$, this extra knowledge retrieved from the contextual KG are integrated into the origin question as a part of the prompt. The prompt is input to the LLM and the LLM perform the reasoning step to determine whether the sufficient information has been retrieved from the contextual KG. If yes, the LLM generates the answer based on the retrieved knowledge and its inherent knowledge. Otherwise, the whole process will iterate by starting the first step with new reasoning paths $P$ and relation context set $C$.
@@ 4 Experiments on KG Completion
  • RQ2: Whether different types of entity contexts contribute to enhancing KG reasoning?
@@
  • RQ4: Can KGR3 improve the inference performance for predicting long-tail entities?
@@ 5.3 Implementation
We use ChatGPT (GPT-3.5-turbo) as the backbone LLM for KGR3 by calling OpenAI API. The maximum token length for the generation is set to 256. In all experiments, we set both width $M$ and depth $D_{max}$ to 3 for beam search. We use 5 shots in KGR3-reasoning prompts for all the datasets.
@@ 5.4 Experimental Results
Table 6: Exact match accuracy of KGR3 using ChatGPT as the backbone models on QALD10-en and WWQ. The prior FT (Fine-tuned) and prompting SOTA include the best-known results: $\alpha$: Santana et al. (2022); $\beta$: Xu et al. (2023); $\theta$: Sun et al. (2024)
Since KGR3 uses external KGs and contextual information to enhance LLM, we first compare it with those methods leveraging external knowledge as well. As we can see in Table 6, even if KGR3 is a training-free prompting-based method and has natural disadvantage in comparison with those fine-tuning methods trained with data for evaluation, KGR3 still achieves new SOTA performance in both datasets. If comparing with other promoting-based methods with ChatGPT as backbone models (especially ToG), KGR3 can win the competition on all datasets.
It is noteworthy that other prompting-based methods rely solely on triple knowledge from KGs, whereas KGR3 allows the LLM to leverage additional contextual information for more precise reasoning on KGs. This is likely the primary reason why KGR3 outperforms other prompting-based methods. To verify this, we evaluated a variant of KGR3 that excludes contextual information for comparison. As shown in Table 6, incorporating contextual information results in a relative increase of 43.6% and 17.1% in Exact Match (EM) on QALD10-en and WWQ, respectively. These experimental results support our hypothesis that KGQA methods can significantly benefit from the integration of contextual information.
@@ 6 Conclusion
This work points out several critical shortcomings of triple-based KGs, including their inability to represent diverse knowledge flexibly and perform complex knowledge reasoning accurately, due to the lack of contextual information. By highlighting these limitations, we underscore the necessity of moving beyond triple-based representation for KGs and introduce the concept of contextual KGs. Contextual KGs integrate rich contextual data, such as temporal, geographic, and provenance information, thus providing a more comprehensive and accurate representation of knowledge. This enhanced representation supports more effective reasoning by leveraging the added layers of contextual information.
To verify the effectiveness of incorporating contextual KGs on KG representation and reasoning, we present KGR3, a novel reasoning paradigm for KGs that integrates LLMs (LLMs) with contextual KGs to address the limitations of traditional triple-based KG reasoning methods. Through extensive experiments on KG completion and KG question answering tasks, we demonstrated that incorporating contextual information significantly improves the performance of existing models. Our results underscore the importance of context in capturing the complexity and richness of real-world knowledge, enabling more nuanced and accurate inferences.
In conclusion, the introduction of contextual KGs represents a significant step forward in the evolution of KGs, offering a more sophisticated and comprehensive approach to knowledge representation and reasoning. This work opens new avenues for future research and applications, highlighting the potential of contextual KGs and LLMs in advancing the field of artificial intelligence.

### v3

# Context Graph

@@ Abstract
Knowledge Graphs (KGs) are foundational structures in many AI applications, representing entities and their interrelations through triples. However, triple-based KGs lack the contextual information of relational knowledge, like temporal dynamics and provenance details, which are crucial for comprehensive knowledge representation and effective reasoning. Instead, Context Graphs (CGs) expand upon the conventional structure by incorporating additional information such as time validity, geographic location, and source provenance. This integration provides a more nuanced and accurate understanding of knowledge, enabling KGs to offer richer insights and support more sophisticated reasoning processes. In this work, we first discuss the inherent limitations of triple-based KGs and introduce the concept of CGs, highlighting their advantages in knowledge representation and reasoning. We then present a context graph reasoning CGR3 paradigm that leverages large language models (LLMs) to retrieve candidate entities and related contexts, rank them based on the retrieved information, and reason whether sufficient information has been obtained to answer a query. Our experimental results demonstrate that CGR3 significantly improves performance on KG completion (KGC) and KG question answering (KGQA) tasks, validating the effectiveness of incorporating contextual information on KG representation and reasoning.
Context Graph
@@ 1 Introduction
While this triple-based structure offers clear semantics and precision through the use of schemas and ontologies, it loses the contextual information of knowledge and falls short in capturing the complexity and richness of real-world knowledge Dong (2023). Since we cannot clearly model the knowledge in a domain only with entities and relations, many recent KGs Pellissier Tanon et al. (2020); Tharani (2021) are designed to be semi-structured: they leverage the clear semantics of structured data provided by the rigidity of schemas (i.e., ontologies) while also embracing the flexibility of unstructured data. Such KGs integrate multi-modal knowledge, including entity description, images, timestamps and other metadata, all of which can be regarded as the contexts of triple knowledge. In this paper, we refer to this type of KGs as contextual graphs (CGs). By incorporating these semantic contexts, CGs provide a more comprehensive and nuanced representation of knowledge, extending beyond the traditional triple-based approach. This enables KGs to possess more advanced capabilities in knowledge representation and reasoning.
Moreover, large language models (LLMs), pre-trained on vast text corpora, have exhibited strong semantic understanding capability Brown et al. (2020a). And the use of LLMs for KG reasoning has become a research hotspot Wei et al. (2023); Liao et al. (2024); Sun et al. (2024). However, KGs may contain numerous entities and relations, but not all entities and relationships are fully annotated and connected, leading to data sparsity. This sparsity results in a lack of sufficient contextual information for the LLM during inference. On the other hand, LLMs are better at handling unstructured data rather than structured triples. Considering that CGs can provide unstructured contextual information for LLM reasoning, the synergy between CGs and LLMs holds significant potential for advancing the field of knowledge reasoning.
In this paper, we will first give a brief discussion on the limitations of the triple-based KGs and give the specific definition of CGs. To validate the effectiveness of contexts on enhancing knowledge representation and reasoning, we propose a novel context graph reasoning paradigm, named CGR3, which leverage the strong reasoning power of LLMs to firstly retrieve candidate entities and related contexts from KG, and rank the candidate entities based on retrieved context, and then reason whether sufficient information is retrieved to answer the question. Experimental results demonstrate that our proposed paradigm CGR3 enhances the performance of existing models on the tasks of KG completion (KGC) and KG question answering (KGQA), which are two of fundamental reasoning tasks over KGs.
@@
  • Point out the limitations of the current triple-based KGs, and introduce the concept of Context Graph, which has more advanced capabilities in knowledge representation and reasoning.
  • Propose a context-enhanced KG reasoning paradigm, CGR3, which leverages the LLM to perform CG reasoning based on related contexts. Experimental results on KGC and KGQA support our intuition that the integration of contextual data can contribute to effective KG reasoning.
2 Context Graph
In this section, we first discuss on the limitations of triple-based KGs, caused by the absence of contextual information. Moreover, we point out the effects of contextual information on knowledge representation and reasoning, then categorize and interpreter different types of contexts in the KGs. Finally, we formally define CGs as well as two knowledge reasoning tasks over CGs.
@@ 2.2 The Effects of Contextual Information
To address the limitations of triple-based KGs, a promising approach is to attach contextual data to factual triples. For instance, several KGs, such as YAGO and the Yahoo Knowledge Graph, include meta-information with their facts, such as the time of validity, the geographic location of a fact, and provenance information. By integrating such data, CGs can offer a more comprehensive and accurate representation of knowledge, thereby enabling more effective reasoning.
@@
During the process of knowledge reasoning, contextual information within CGs can be leveraged to associate entities that are not directly connected by identifying similar contexts. This capability is particularly useful for making connections and drawing inferences that go beyond the predefined relation set of a triple-based KG. Moreover, contextual information provides additional knowledge, allowing for larger knowledge coverage and greater flexibility compared to triple-based KGs. Specifically, contextual information can be used to answer complex reasoning questions, such as those involving qualifiers or specific conditions that are often hidden within contextual data. For instance, answering a question about "which company is Apple’s biggest competitor in the global smartphone market" would require integrating quantitative data, temporal information, and detailed market dynamics analysis with basic entity and relation information in KGs, as shown in Figure 1(d). CGs thus enable the handling of such intricate queries by providing a richer and more detailed knowledge base.
2.3 Categories of Contextual Data
@@
As shown in Figure 2, contextual data can be roughly classified into two categories, i.e., entity contexts and triple contexts.
@@ 2.4 Problem Specification
A Context Graph (denoted as $\mathcal{CG}=\{\mathcal{E},\mathcal{R},\mathcal{Q},\mathcal{EC},\mathcal{RC}\}$) can be represented as a set of factual quadruples in the form of $(h,r,t,rc)\in\mathcal{Q}$, where $h,t\in\mathcal{E}$, $r\in\mathcal{R}$ and $rc\in\mathcal{RC}$. The notations $h$ and $t$ denote the head and the tail entity of a factual quadruple, $r$ denotes the relations between $h$ and $t$, and $rc$ denotes . $\mathcal{EC},\mathcal{RC}$ are the set of entity contexts and relation contexts. Each entity $e\in\mathcal{E}$ and its entity context $ec\in\mathcal{EC}$ form a complete entity representation $(e,ec)$.
@@
Given a natural question $nq$ and its topic entity $e_{topic}\in\mathcal{E}$, KGQA aims to retrieve related knowledge by generating structured queries or sampling subgraphs from $\mathcal{KG}$ and predict the answer $a$ based on retrieved knowledge, i.e, $a=f(nq,e_{topic},\mathcal{KG})$. For performing QA over a contextual KG, the prediction function can be reformulated as $f(nq,e_{topic},\mathcal{CG})$.
Figure 3: The pipeline of the CGR3 paradigm.
@@ 3 Methods
In this section, we introduce CGR3, a novel context graph reasoning paradigm that leverages LLMs to perform knowledge reasoning tasks based on structured and contextual semantics. We aim to utilize the complementary relationship between both semantics to improve the reliability and explainability of the reasoning process.
As shown in For triple-based KGs, we begin by augmenting the KG with necessary contextual information extracted from relevant databases, a step that can be omitted if the KG is already a CG. The CGR3 paradigm consists of three main steps: The Retrieval step is to retrieve candidate entities and related contexts from the CG based on the given question; the Ranking step involves ranking candidate entities based on the contexts and the given question; the Reasoning step is to exploit the LLM to determine whether sufficient information is retrieved. If sufficient information is available, the answer will be generated. If not, the whole processes iterates by retrieving new information based on the top-ranked candidate entities. We give a detailed description of the proposed context-aware paradigms for the KGC and the KBQA tasks
@@ 3.1 Context Extraction
Currently, commonly used KG datasets, such as FB15k237, YAGO3-10, and Wikidata5M, are encyclopedic KGs that encapsulate general knowledge about the real world. These KGs are typically developed by domain experts by applying named entity recognition and relation extraction techniques on Wikipedia. However, during this construction process, the rich contexts surrounding the entities are often omitted. Recent studies Wang et al. (2021, 2022b) have proposed to incorporate entity labels and descriptions as supplementary information for KGs. Nevertheless, the labels and descriptions are insufficient to replace the specific contexts associated with KG triples, thereby limiting their effectiveness in addressing diverse knowledge reasoning problems.
@@ 3.2 Knowledge Graph Completion
In this section, we demonstrate a new context-enriched KGC method based on our proposed CGR3 paradigm. Since KGC can be considered as an entity ranking task for single-hop reasoning questions, it is not necessary to perform iterative reasoning processes. Thus, the reasoning step is omitted for this task.
@@ 3.3.3 Step 3: Context-aware Reasonin
Upon obtaining the new top-$M$ reasoning paths $P$ and updating relation context list $C$, this extra knowledge retrieved from the CG are integrated into the origin question as a part of the prompt. The prompt is input to the LLM and the LLM perform the reasoning step to determine whether the sufficient information has been retrieved from the CG. If yes, the LLM generates the answer based on the retrieved knowledge and its inherent knowledge. Otherwise, the whole process will iterate by starting the first step with new reasoning paths $P$ and relation context set $C$.
@@ 4 Experiments on KG Completion
  • RQ2: Whether different types of entity contexts contribute to enhancing knowledge reasoning?
@@
  • RQ4: Can CGR3 improve the inference performance for predicting long-tail entities?
@@ 5.3 Implementation
We use ChatGPT (GPT-3.5-turbo) as the backbone LLM for CGR3 by calling OpenAI API. The maximum token length for the generation is set to 256. In all experiments, we set both width $M$ and depth $D_{max}$ to 3 for beam search. We use 5 shots in CGR3-reasoning prompts for all the datasets.
@@ 5.4 Experimental Results
Table 6: Exact match accuracy of CGR3 using ChatGPT as the backbone models on QALD10-en and WWQ. The prior FT (Fine-tuned) and prompting SOTA include the best-known results: $\alpha$: Santana et al. (2022); $\beta$: Xu et al. (2023); $\theta$: Sun et al. (2024)
Since CGR3 uses external KGs and contextual information to enhance LLM, we first compare it with those methods leveraging external knowledge as well. As we can see in Table 6, even if CGR3 is a training-free prompting-based method and has natural disadvantage in comparison with those fine-tuning methods trained with data for evaluation, CGR3 still achieves new SOTA performance in both datasets. If comparing with other promoting-based methods with ChatGPT as backbone models (especially ToG), CGR3 can win the competition on all datasets.
It is noteworthy that other prompting-based methods rely solely on triple knowledge from KGs, whereas CGR3 allows the LLM to leverage additional contextual information for more precise reasoning on KGs. This is likely the primary reason why CGR3 outperforms other prompting-based methods. To verify this, we evaluated a variant of CGR3 that excludes contextual information for comparison. As shown in Table 6, incorporating contextual information results in a relative increase of 43.6% and 17.1% in Exact Match (EM) on QALD10-en and WWQ, respectively. These experimental results support our hypothesis that KGQA methods can significantly benefit from the integration of contextual information.
@@ 6 Conclusion
This work points out several critical shortcomings of triple-based KGs, including their inability to represent diverse knowledge flexibly and perform complex knowledge reasoning accurately, due to the lack of contextual information. By highlighting these limitations, we underscore the necessity of moving beyond triple-based representation for KGs and introduce the concept of CGs. CGs integrate rich contextual data, such as temporal, geographic, and provenance information, thus providing a more comprehensive and accurate representation of knowledge. This enhanced representation supports more effective reasoning by leveraging the added layers of contextual information.
To verify the effectiveness of incorporating contexts on knowledge representation and reasoning, we present CGR3, a novel knowledge reasoning paradigm that integrates LLMs (LLMs) with CGs to address the limitations of traditional triple-based knowledge reasoning methods. Through extensive experiments on KG completion and KG question answering tasks, we demonstrated that incorporating contextual information significantly improves the performance of existing models. Our results underscore the importance of context in capturing the complexity and richness of real-world knowledge, enabling more nuanced and accurate inferences.
In conclusion, the introduction of CGs represents a significant step forward in the evolution of KGs, offering a more sophisticated and comprehensive approach to knowledge representation and reasoning. This work opens new avenues for future research and applications, highlighting the potential of CGs and LLMs in advancing the field of artificial intelligence.